Show how you would solve the following systems of equations: x + Y + Z = 6 2X – Y + 3Z = 8 3X – 2Y – Z = -17
Answer by Alan3354(69443)

Show how you would solve the following systems of equations:
x + Y + Z = 6
2X – Y + 3Z = 8
3X – 2Y – Z = -17
Eliminate y:
Add eqns 1 & 2
x + Y + Z = 6
2X – Y + 3Z = 8
---------------- Add
3x + 4z = 14
--------
Multiply eqn 1 by 2, and add to eqn 3
2x + 2y + 2z = 12
3X – 2Y – Z = -17
------------------ Add
5x + z = 5
3x +4z = 14
----------- Now it's 2 eqns in x & z. Can you finish it?